### v1

# Molecular Mechanisms Underlying the Effects of Urea and the Structural Dynamics of Bovine Serum Albumin

Abstract
The disruption of protein structures by denaturants like urea is well-studied, though its molecular mechanisms remain unclear. Using Molecular Dynamics (MD) simulations, we investigated how urea affects the structural stability of Bovine Serum Albumin (BSA) at concentrations from 0 M to 5 M. Our results reveal that urea induces a dehydration-rehydration cycle, characterized by displacement and partial replacement of water molecules in BSA’s hydration shell. At low concentrations, urea reduces protein-water hydrogen bonds while enhancing protein-urea interactions. At higher concentrations, urea aggregation limits these interactions, promoting rehydration and changes in tertiary structure, while secondary structure remains largely intact. These findings provide insights into the mechanisms of protein denaturation and stability by urea.
preprint: AIP/123-QED
I Introduction
The study of urea and other denaturing agents in aqueous protein solutions has been a central topic in biochemistry for decades. The ability of these compounds to alter the three-dimensional structure of proteins is well documented; however, the exact molecular mechanism underlying this process remains a subject of debate Wallqvist, Covell, and Thirumalai (1998); Frank and Franks (1968); Åstrand et al. (1991); Rezus and Bakker (2006); Soper, Castner Jr, and Luzar (2003); Gahtori, Gunwant, and Pandey (2023); Stumpe and Grubmüller (2007); Lindgren and Westlund (2010); Moeser and Horinek (2014); Matubayasi (2018); Nnyigide, Lee, and Hyun (2018); Kumaran and Ramamurthy (2011); Ma, Pazos, and Gai (2014); Sinibaldi et al. (2008); Monhemi et al. (2014).
Several mechanisms have been proposed to explain urea-induced protein denaturation. Traditionally, two main models have been considered: the direct mechanism, where urea interacts with stabilizing intramolecular forces via hydrogen bonding and van der Waals interactions; and the indirect mechanism, where urea alters solvent properties, thereby weakening hydrophobic interactions Rossky (2008); Das and Mukhopadhyay (2009); Niether et al. (2018); Hua et al. (2008); Jha and Marqusee (2014); Guckeisen, Hosseinpour, and Peukert (2021). Recent studies, however, suggest that these mechanisms are not mutually exclusive but may act synergistically, depending on the experimental conditions Khan et al. (2019); Paladino et al. (2022); Lim, Rösgen, and Englander (2009).
The impact of urea on protein stability is influenced by multiple factors, including denaturant concentration, temperature, pH, and thermodynamic parameters Miller and Levinger (2022). Comparisons between urea and guanidinium chloride have revealed significant differences in their effects on protein conformation Baptista et al. (2023); Candotti et al. (2013); Maity et al. (2020). Computational studies indicate that urea exhibits a particular affinity for hydrophobic residues, facilitating solvation and disrupting the hydrophobic collapse Lapidus et al. (2007), ultimately leading to loss of native structural integrity Candotti et al. (2013). Recent molecular dynamics simulations have further demonstrated that, under acidic conditions, urea undergoes a localized reorientation near protonated acidic residues (Asp/Glu), acting predominantly as a hydrogen bond acceptor. However, these changes do not significantly affect urea’s interactions with the protein backbone or nonpolar residues, supporting a direct and consistent denaturation mechanism across pH conditions de Oliveira and Martínez (2020). Complementing these computational findings, surface-sensitive vibrational spectroscopy studies have experimentally confirmed that urea interacts directly with hydrophobic side chains at protein-like interfaces. This interaction reduces the hydrophobic barrier, promotes interfacial water ordering, and disrupts side chain organization, further reinforcing the direct role of urea in destabilizing the protein hydrophobic core Gahtori, Gunwant, and Pandey (2023).
Additionally, molecular dynamics (MD) simulations have shown that the dipolar orientation of urea in aqueous solution can temporarily stabilize certain secondary structural elements before their complete disruption Baptista et al. (2023). This finding suggests that denaturation is not an instantaneous event but a dynamic process governed by the competition between protein-protein and protein-solvent interactions Paladino, Vitagliano, and Graziano (2023); Gooran and Kopra (2024).
In this study, we employ MD simulations to investigate the effects of varying urea concentrations on bovine serum albumin (BSA) structure. This protein was selected due to its ability to undergo reversible conformational changes in response to pH variations Scanavachi et al. (2020). Specifically, we focus on the N-isoform of BSA at pH 6.5, a physiologically stable and functional conformation. Our results exhibit strong agreement with both experimental and previous computational studies, reinforcing the validity of our approach and contributing to a deeper understanding of urea-induced protein denaturation.
II Materials and methods: Systems and simulations
We investigated the impact of urea on the structural integrity of bovine serum albumin (BSA) in aqueous solutions across varying urea concentrations. The computational model utilized for urea was the Boek model Boek, Briels, and Feil (1994) with nonbonded interactions and covalent parameters sourced from the GROMOS 54A7 force field Schmid et al. (2011). For water, we employed the Simple Point Charge Extended (SPC/E) model Berendsen, Grigera, and Straatsma (1987), as developed by Berendsen and colleagues. The initial atomic coordinates for the BSA model were obtained from the crystal structure of bovine serum albumin deposited in the Protein Data Bank (PDB code 4F5S) Bujacz (2012). The topology of BSA at a standard pH of 6.5 was obtained from previous work and validated through Small-Angle X-ray Scattering (SAXS) and MD simulations Scanavachi et al. (2020). The net charge of the protein under these conditions was +11, and sodium (Na+) counterions were added to neutralize the system.
Intermolecular interactions were explicitly accounted for using the GROMACS 2020.6 molecular dynamics simulation package Spoel et al. . As a reference, a control system devoid of urea was first simulated, with water molecules randomly distributed around the protein (see Control System in Table 1).
Systems
Water
molecules
Contraions
Na+
Urea
molecules
Urea
concentrations
Control 61,212 11 0 0 M
1M 57,312 11 1,099 1 M
2M 53,356 11 2,198 2 M
3M 49,235 11 3,278 3 M
4M 44,840 11 4,396 4 M
5M 40,963 11 5,468 5 M
Table 1: The simulated systems. Bovine Serum Albumin (BSA) in water (control), and BSA in urea-water solutions at varying urea concentrations ranging from 1 to 5 M.
[…]
After optimization of the control system, the final equilibrated conformation was used as the reference structure for subsequent simulations involving urea. Urea molecules were randomly inserted into this initial configuration (Figure 1), yielding five simulation boxes containing BSA solvated in water–urea mixtures at increasing concentrations ranging from 1 M to 5 M (see Table 1). Each system was subjected to the same two-step optimization protocol previously applied to the control system, ensuring methodological consistency across all conditions. Production simulations were then carried out for 500 ns. From the final 8 ns of each trajectory, eight distinct configurations were randomly selected to serve as starting points for additional 10 ns simulations. Each replicate was initiated with a different set of initial velocities to enhance statistical sampling. This protocol enabled a detailed investigation of the long-timescale dynamics and intermolecular interactions within the protein–urea–water systems.
Figure 1: Water box simulation featuring urea, sodium, and Bovine Serum Albumin (BSA). BSA divided into three domains: D1 (yellow, residues 1-185), D2 (violet, residues 186-378), and D3 (light blue, residues 379–576). Sodium ions (Na+) depicted in green. Urea molecules are represented by cyan carbon, red oxygen, blue nitrogen, and white hydrogen atoms.
The optimization and production processes of both the control system and BSA in water-urea solutions were simulated using a cubic box with periodic boundary conditions, at $300.15$ K and $1$ bar of pressure, employing a velocity rescale thermostat Bussi, Donadio, and Parrinello (2007) and Berendsen barostat Berendsen et al. (1984), respectively. The coupling time constants utilized in the thermostat and barostat were $0.1$ ps and $1.0$ ps, respectively. Electrostatic interactions were computed using the PME method Abraham and Gready (2011) with a cutoff for Van der Waals and Coulomb interactions set at $1.0$ nm. The LINCS algorithm Hess et al. (1997) was employed to constrain solute bonds. A time step of $2$ fs was maintained throughout the simulations.
Hydrogen Bonds. The number of hydrogen bonds per molecule was calculated using the standard GROMACS utility gmx hbond. A hydrogen bond was considered to exist when the donor–acceptor distance was $\leq$0.35 nm and the acceptor–donor–hydrogen angle (A–D–H) was $\leq$30, by commonly accepted geometric criteria in GROMACS.
III Results and Discussion
All analyses are based on the average of eight independent simulations, each 10 ns in duration, conducted for each system studied (control without urea and urea:water mixtures from 1 to 5 M). Before data collection, all systems underwent 500 ns of hydration/solvation to ensure stabilization.
III.1 Hydrogen Bond Analysis
Initially, we analyzed the formation of protein-water hydrogen bonds (HBs) as a function of increasing urea concentration. For each of the eight systems corresponding to each concentration, the calculations were averaged over the final 8 ns of the simulation.
As shown in Figure 2, the presence of urea in the BSA–water system leads to a marked reduction in protein–water HBs. At 1 M urea, a decrease of approximately 25% is observed, which increases to nearly 45% as the concentration increases from 2 M to 5 M. Interestingly, beyond this point, the number of protein–water HBs plateaus at around 700, indicating that urea-induced dehydration reaches a threshold beyond which further displacement of water molecules from the protein surface is limited.
Figure 2: Total number of hydrogen bonds (HBs) formed between BSA and the components of urea:water mixtures as a function of urea concentration. Light blue bars represent the average number of HBs between the protein and water molecules, while light green bars correspond to HBs formed between the protein and urea molecules. The horizontal orange line indicates the average number of protein–water HBs in the absence of urea (0 M), used as a reference. Error bars represent the standard deviation across eight replicates. The data illustrate the progressive substitution of protein–water HBs by protein–urea HBs with increasing urea concentration, while the total number of protein–solvent HBs remains relatively constant. The supplementary material includes figures showing the HBs distributions for protein–water, protein–urea, and urea–urea interactions, normalized by the number of molecules of each species.
Similarly, Figure 2 displays the average number of hydrogen bonds between the protein and urea molecules, highlighted in light green. Notably, the decrease in protein-water HBs with increasing urea concentration is proportional to the increase in protein-urea HBs. This observation suggests that as urea concentration rises, the protein undergoes progressive dehydration, accompanied by a corresponding increase in urea solvation. In this case, we observe that as the concentration increases, the average number of HBs between urea molecules and the protein initially rises but eventually stabilizes around 600. This plateau suggests that both urea-induced dehydration and urea solvation may reach a threshold, beyond which additional urea molecules do not significantly alter the hydration state of the protein.
Taking the average number of protein–water HBs at 0 M urea as a reference, Figure 2 shows that BSA consistently maintains over 1,200 hydrogen bonds across all urea concentrations. Although increasing urea concentration induces protein dehydration, resulting in the loss of protein–water HBs, a compensatory gain in protein–urea HBs is observed. However, the overall increase in the total number of HBs formed between BSA and the urea:water binary mixtures is approximately 2%. This suggests that urea molecules increasingly interact with the BSA, potentially displacing water molecules and altering the hydrogen bonding network around the protein.
In addition, as the urea concentration increases, the total number of HBs between urea molecules grows monotonically (see Figure 3a in the supplementary material). However, when normalized by the number of urea molecules in the system (see Figure 3b in the supplementary material), the number of HBs per urea molecule exhibits an asymptotic increase with concentration. Notably, beyond 3 M, this value begins to stabilize, whereas at 1 M it remains relatively low. This behavior suggests that as the concentration increases, urea molecules form a more interconnected hydrogen-bond network that gradually approaches saturation.
III.2 Radial Distribution Function (RDF)
To support our initial findings, we analyzed the radial distribution functions (RDFs) calculated between the center of mass (COM) of the protein and that of water molecules, and separately, between the protein COM and the COM of urea molecules.
In Figure 3(a), the RDF between the COM of the protein and that of water molecules shows that, even in the presence of urea, water molecules can approach the protein COM as closely as $\sim$0.20 nm. This suggests that local hydration near solvent-accessible regions of the protein core is maintained across all urea concentrations. Additionally, a secondary peak appears around 0.28–0.30 nm, which corresponds to the outer hydration shell. This feature is most prominent in the absence of urea (0 M) and progressively decreases in intensity with increasing urea concentration, indicating a disruption of long-range water structuring around the protein. However, a deviation from this trend is observed at 2M urea, suggesting a shift in molecular behavior. This deviation is difficult to correlate directly with the total average number of HBs between the protein and water molecules (Figure 2), as the values are not normalized with respect to the number of solvent molecules.
Figure 3: Radial distribution functions (RDFs) calculated between the center of mass (COM) of the protein and the COM of solvent components at different urea concentrations. (a) RDF between the protein COM and the COM of water molecules for the range 0.1-1.0 nm. (b) RDF between the protein COM and the COM of urea molecules for the same distance range. A redistribution of solvent density near the protein is observed as urea concentration increases.
By analyzing the normalized average number of hydrogen bonds between protein and water (Figure 1 in the Supplementary Material) and its variation in the presence of urea, we show that the result is consistent with the observations in 4(a). From this, it can be observed that from 0 M to 2 M, the normalized average number of HBs between the protein and water decreases, reaching a minimum at 2 M. Beyond this minimum, the normalized average number of HBs between the protein and water increases, accompanied by the corresponding rehydration of the protein. Although the number of water molecules decreases by approximately 30% when comparing 1 M and 5 M urea concentrations (Figure 2), the normalized average number of HBs remains almost the same at both concentrations.
When analyzing the RDF between urea molecules and the protein, shown in Figure 3(b), two distinct peaks are evident at 1 M urea. The first peak ($\sim$0.20–0.30 nm) corresponds to urea molecules close to the protein surface, indicating short-range interactions. The second peak ($\sim$0.40–0.50 nm) is associated with urea molecules located in the outer solvation shell. As the urea concentration increases, the intensity of both peaks progressively decreases, suggesting a reduced presence of urea near the protein and a possible redistribution into the bulk solvent.
The analysis based on the number of HBs between urea molecules and the protein, normalized by the number of urea molecules (Fig. 2 in the supplementary material), reinforces and complements the observations in Fig. 3(b). As the urea concentration increases, the number of HBs formed with the protein remains relatively constant at high concentrations, leading to a monotonic decrease in the average normalized HB count.
We interpret the sustained HBs count at high urea concentrations as evidence that protein solvation by water molecules persists, while direct interactions with urea molecules diminish. Complementarily, we analyzed the formation of hydrogen bonds between the protein main chain, protein side chain, and water and urea molecules (see Figures 4 and 5 in the supplementary material).These results suggest that increasing urea concentration does not induce significant changes in the interactions between urea or water molecules and the protein main chain, but does markedly affect their interactions with the protein side chains. This supports the notion that urea–protein interactions are more likely to occur with the side chains rather than the backbone.
When examining the radial distribution function (RDF) between urea carbons to better understand urea’s role in its interactions with the protein (Figure 4). At lower urea concentrations (1M to 2M), there is an increase in both the central and second peaks of the RDF. At higher concentrations ($>$2M), however, a decrease in the central peak is observed, indicating a shift in behavior compared to the trends at 1M and 2 M. This decrease at concentrations above 2 M can be attributed to specific geometric arrangements that urea molecules adopt during aggregation, as shown in Figure 4.
Figure 4: RDF between the carbon atoms of urea molecules and different snapshots of the final configurations at each urea concentration.
The inset images in Figure 4 reveal these changes in urea aggregation geometry, which are consistent with the model proposed by Susan and collaborators Atahar et al. (2019). The complementary analysis made with the normalized hydrogen bond distribution (see Fig. 3b in the supplementary material) among urea molecules aligns with the aggregation changes depicted in Fig. 4. This suggests that as urea concentration increases, the preference for urea-urea interactions also increases, and this tendency toward self-interaction appears to become dominant at higher concentrations.
III.3 Radius of Gyration (RG), Root Mean Square Deviation (RMSD) and Solvent Accessible Surface Area (SASA)
All the parameters presented in this section were analyzed over the last eight nanoseconds of the eight replica simulations performed at each urea concentration. Reported values correspond to averages computed across these datasets.
We analyzed the protein’s radius of gyration as a function of urea concentration and observed that the average radius increases with increasing urea concentration. At 2 M, the radius of gyration increases more pronouncedly, as shown in Figure 5(a). At this concentration, the interaction between the protein and water is minimal, corresponding to the lowest number of hydrogen bonds per water molecule formed between the protein and water molecules (Figure 1 in the supplementary material). Overall, the presence of urea increases the average radius of gyration compared to the system without urea. However, these changes remain modest, not exceeding 6%, suggesting that significant conformational alterations are unlikely. This finding is consistent with the observations of Yasuda et al. Hayashi et al. (2007), who reported that within this range of urea concentrations and temperatures, the protein structure remains relatively stable, with only minor perturbations driven by solvent dynamics.
Figure 5: (a) Mean Radius of gyration vs urea concentration. (b) Mean of Root Mean Square Deviation vs urea concentration. (c) Mean solvent accessible surface area of the protein vs urea concentration.
The analysis of the average RMSD as a function of urea concentration, presented in Figure 5(b), was performed by comparing the structures concerning the one obtained from the Protein Data Bank (PDB code: 4F5S).
Most urea concentrations reduce the average RMSD compared to the protein’s native state in the absence of urea, suggesting structural stabilization. Notably, RMSD values at 1M and 3M urea are comparable, whereas significantly lower values at 4M and 5M indicate enhanced stabilization. In contrast, at 2M urea, the average RMSD increases relative to 0M, suggesting structural deviation. The destabilization observed at 2M urea correlates with changes in the protein’s radius of gyration under the same conditions. These findings suggest a relationship between structural destabilization and protein expansion at this concentration. Moreover, the data presented in this figure support the RG analysis, reinforcing the conclusions regarding the role of water and its interaction with the protein.
When analyzing the mean value of the total solvent-accessible surface area ( Figure 5(c)), it is evident that the introduction of urea leads to an increase in solvent exposure. In the absence of urea, the protein exhibits the lowest SASA, which increases monotonically up to a concentration of 4 M. The percentage increases in the average SASA values slightly exceed 9% only at a urea concentration of 4 M; in all other cases, they remain below this level compared to the protein in the absence of urea. Interestingly, at higher urea concentrations, deviations from this monotonic behavior emerge. While these deviations do not contradict the overall trend, they underscore the need for a more detailed analysis, particularly one that focuses on how the SASA changes across individual protein domains. Given the structural complexity of proteins such as BSA, a domain-specific SASA analysis is essential for understanding how different regions of the protein respond to varying urea concentrations, especially regarding solvent exposure and potential local conformational changes.
To refine the analysis, we investigated the SASA by domain and the inter-domain distances to identify the protein regions most affected by hydration changes or variations in urea concentration (Figure 6).
Across all urea concentrations, Domain 1 consistently exhibits the lowest solvent-accessible surface area. In contrast, Domain 2, consistent with previous studies Scanavachi et al. (2020), demonstrates the greatest propensity for surface exposure, displaying the highest SASA values at nearly all urea concentrations. Notably, for all concentrations, the solvent-exposed surface area of all three domains increases compared to the system without urea.
Figure 6: (a) Mean solvent accessible surface area per domain vs urea concentration. (b) Mean distance between the center of mass of the domains as a function of concentration. The errors in the mean values shown in the figures are small, making them barely visible without enlarging the figure.
The presence of urea increases the SASA of each domain, observing that the contribution at total SASA is different, and dependent on urea concentration. To better understand the role of each domain, we analyzed the interdomain distance in Figure 6(b).
As an initial observation, Figure 6(b) shows that at 0M urea, the distances between D1-D3 and D2-D3 are approximately equal, both near 3.6 nm. In contrast, the D1-D2 distance is around 3.2 nm, suggesting that the centers of mass of the domains form an isosceles triangle. As the urea concentration increases, the distance between the D1 and D3 domains increases relative to the control system without urea. However, the distances between the D1-D2 and D2-D3 domains fluctuate around the reference value with the addition of urea. Interestingly, at 2 M urea, the inter-domain distances D1-D3 tend to increase, followed by a decrease at 3 M. Additionally, at 4 M and 5 M urea, the D1-D2 distance exhibits an opposite behavior compared to the D1-D3 and D2-D3 distances. This suggests that Domain 3 is the most affected by urea concentration, as it moves away from Domain 1 and 2, leading to a greater solvent-exposed surface area for this region (Figure 6(a).
III.4 Secondary Structure
To gain deeper insights into how urea impacts the protein structurally, we analyzed its secondary structure as a function of urea concentration.
As shown in Figure 7(a), the presence of urea induces a slight increase in the number of amino acids participating in the secondary structure. This increase, however, is modest, with the largest observed difference being just over 4% when comparing the absence of urea to 5M. Overall, urea concentration does not appear to play a significant role in driving changes to the secondary structure. The variations in the number of amino acids involved in the secondary structure between 1M and 5M are minimal, amounting to slightly more than 1%.
Figure 7: Mean number of amino acids (AA) involved in different types of secondary structure as a function of urea concentration. (a) Total number of amino acids forming any type of secondary structure. (b) Number of amino acids forming $\alpha$-helices. (c) Number of amino acids forming turns.
A more detailed secondary structure analysis reveals a clear trend of increasing $\alpha$-helix formation with urea concentration (Figure 7(b)). At all concentrations, the number of $\alpha$-helices is consistently higher than at 0M, with the largest increase observed at 1M. Beyond 2M, the number of amino acids forming $\alpha$-helices increases almost monotonically with concentration.
In contrast, Figure 7(c) shows that the average number of amino acids forming turns is generally reduced compared to 0 M, except at 2 M, where it peaks above the initial value. Beyond 2 M, the number of turns decreases monotonically with increasing urea concentration.
These findings suggest urea reinforces the secondary structure by promoting $\alpha$-helix formation. However, this is accompanied by a general, albeit smaller, reduction in the number of amino acids forming turns. As a preliminary hypothesis, it can be proposed that the structural changes induced by urea primarily affect the protein’s tertiary structure. The observed difference of approximately 3-4% in the average number of amino acids in the secondary structure corresponds to around fifteen residues. This numerical difference implies that the loss of hydrogen bonds observed in the protein is not directly associated with reductions in its secondary structure.
IV Conclusion
This study employed molecular dynamics simulations to explore the effects of urea on the structural behavior of bovine serum albumin (BSA) at physiological pH. Across the range of urea concentrations tested (0–5 M), our results reveal that urea primarily induces changes at the tertiary structure level, with the secondary structure remaining largely conserved, particularly the $\alpha$-helical content. This structural integrity is consistent with previous experimental reports and highlights the stability of the N-isoform of BSA under mildly denaturing conditions Murayama and Tomida (2004); Abrosimova, Shulenina, and Paston (2016).
A key observation is the progressive dehydration of the protein surface as urea concentration increases, driven by the replacement of protein–water hydrogen bonds with protein–urea interactions. Despite this shift, the total number of hydrogen bonds involving the protein remains nearly constant, indicating a dynamic equilibrium in solvation.
Radial distribution function analysis and solvation metrics (RDF, SASA, Rg) further support a modest expansion of the protein, especially at intermediate urea concentrations, with domain-specific differences in solvent accessibility, most notably in Domain 2. This expansion correlates with subtle structural rearrangements and may reflect early stages of tertiary unfolding.
At higher concentrations, urea–urea interactions become increasingly dominant, leading to a saturation of protein–urea contacts and promoting partial rehydration of the protein. This saturation arises when all accessible interaction sites on the protein surface are occupied, preventing additional urea molecules from establishing new contacts. As a result, excess urea tends to self-associate Atahar et al. (2019), forming structured aggregates rather than interacting directly with the protein. The system thus reaches a dynamic equilibrium in which the number of protein–urea interactions no longer increases, despite the higher concentration. These findings suggest a complex, concentration-dependent interplay between dehydration, solvation, and domain flexibility, reinforcing the idea that urea-induced denaturation is a gradual and structurally selective process.
While the low number of Na+ ions in the present system limits our ability to evaluate their impact on protein destabilization, future studies could investigate how urea modulates electrostatic interactions under conditions where ionic effects are more pronounced. Previous observations in SDS micellar systems showed that urea can disrupt structure by solvating counterions and weakening electrostatic stabilization Espinosa, Grigera, and Ferrara (2018). A similar mechanism may operate in protein systems, particularly under conditions where ion–protein interactions contribute more significantly to structural integrity. Exploring these effects could provide deeper insight into the interplay between ion binding, solvation dynamics, and urea-induced denaturation.
ACKNOWLEDGMENTS
We wish to thank and fondly remember Dr. Raul Grigera, who taught and shaped us in this beautiful profession, and thank the anonymous reviewers who provided thoughtful feedback on previous drafts of this article. This work was supported by grants from Consejo Nacional de Investigaciones Científicas y Tecnicas (CONICET), Agencia Nacional de Promocion Científica y Tecnologica (ANPCyT) and Universidad Nacional Arturo Jauretche (UNAJ, Argentina).
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
Ethics Approval
Ethics approval is not required.
DATA AVAILBILITY
The data that support the findings of this study are available from the corresponding author upon reasonable request.